please help me solve this equation: Solve for x: 25^x-2=5^3x.
Answer by kapilsinghi(68)

25^x-2=5^3x
25^(x-2)=5^3x
5^2(x-2)=5^3x
2x-4 = 3x
x=-4
kapilsinghi123@gmail.com